1- Who wondered which planet sung soprano?

A) Newton

B) Kepler

C) Plato

2- which of the following is a subjective term?

A) intensity

B) Pitch

C) Frequency

3- Doppler effect occurs when a source of sound moves

A) away from you

B ) toward you

C) both of A&B

D) None of these

4- You are driving along a highway at 35.0 m/s when you hear the siren of a police car approaching you from behind and you perceive the frequency as 1310 Hz. You are relieved that he is in pursuit of a different speeder when he continues past you, but now you perceive the frequency as 1240 Hz. What is the frequency of the siren in the police car? The speed of sound in air is 343 m/s.

A)

1360 Hz

B)

1320 Hz

C)

1270 Hz

D)

1300 Hz

5- An elephant can hear sound with a frequency of 15.0 Hz. What is the wavelength of this wave? The speed of sound in air is 343 m/s.
Answer #1

1) (B) Kepler

2) (B) Pitch

3) (C) Both A and B

4) (C) 1270 Hz

fd=f(v-vr)/(v-vs)

1310=f(343-35)/(343-vs)

1310(343-vs)=308f

343-vs=0.235f----(1)

fd=f(v+vr)/(v+vs)

1240=f(343+35)/(343+vs)

1240(343+vs)=378f

343+vs=0.305f----(2)

Add: (1)+(2)

686=0.540f

f= 1270[Hz]